OpenStudy (anonymous):

Fan and Medal for best answer! Ease show how you got it. Solve by Completing the Square. x^2+x=30

OpenStudy (mathmale):

I'll begin by asking you to identify the coefficients of the x^2 and x terms. To complete the square, take HALF of the coeff. of the x term. Square this result. Add this square to both sides of the given equation. Re-write the left side as the square of a binomail. simplify the right side. Find the square root of both sides, being careful to preface the right side with "plus or minus." What next? Suggest that you check your answers thru substituting them into the original equation.

OpenStudy (anonymous):

\[\left( \frac{ 1 }{ 2 } \right) ^2 \]

OpenStudy (mathmale):

Very nice. what is the square of 1/2?

OpenStudy (anonymous):

1 ?

OpenStudy (mathmale):

\[\frac{ 1^2 }{ 2^2 }=??\]

OpenStudy (anonymous):

1 over 4

OpenStudy (mathmale):

Right, or "one fourth." Your equation is\[x^2 + x = 30\] Please add 1/4 to both sides of this equation.

OpenStudy (mathmale):